OpenStudy (anonymous):

The sum of two numbers is 86. One of the numbers is five less than three times the other. What are the two numbers?

OpenStudy (anonymous):

could you show me the equation?

OpenStudy (anonymous):

x+y=86 x=3y-5 3y-5+y=86

OpenStudy (anonymous):

Let the second number be x, then the first number is 3x-5 (three times the number,5 less than that) now the two add up to 86 so x+3x-5=86 x=91/4=23.75 other number is 3x-5 = 66.25